# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ III-C Linear Complexity TGS
$TGS^{+}$ also requires the initial coordinate distribution $\rho(\cdot|\gamma_{\text{+}}^{(0)})$ as an input, which can be pre-computed via the initialization routine described in Algorithm 3. The recursive construct (lines 3-3, 3-3) reduces the $\mathcal{O}(P^{2}M)$ complexity [30] to an $\mathcal{O}(PM)$ complexity. Specifically, for each $i\in\{1$:$P\}$ computing $\widetilde{\pi}_{i}(\gamma_{\text{+}}^{(0)}(\ell_{i})|\gamma_{\text{+}}^{(0)}(\ell_{\bar{i}}))$, $\nu_{i}^{(1)}(\gamma_{\text{+}}^{(0)})$, $\nu_{i}^{(\beta)}(\gamma_{\text{+}}^{(0)})$ and $\widetilde{\rho}(i|\gamma_{\text{+}}^{(0)})$ incurs $\mathcal{O}(M)$ complexity.
Input : $\gamma_{\text{+}}^{(0)}$, $T$, $\alpha$, $\beta$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$, $\rho(\cdot|\gamma_{\text{+}}^{(0)})$, $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$, $[\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$, $[\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$
Output : $[\gamma_{\text{+}}^{(t)}]_{t\text{=}1}^{T}$
1   $P\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\texttt{=}1}^{P},\textsf{\footnotesize row})$;   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\texttt{=}1}^{P},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3       $n\sim\textsf{\small Categorical}([1\mathnormal{:}P],\rho(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}))$;
4       $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))\leftarrow\frac{\alpha\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))}{\nu_{n}^{(1)}\!(\gamma_{\text{+}}^{(t\text{-}1)})}+\frac{(1\text{-}\alpha)\widetilde{\pi}_{n}^{\beta}\!(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))}{\nu_{n}^{(\beta)}\!(\gamma_{\text{+}}^{(t\text{-}1)})}$;
5       $\gamma_{\text{+}}^{(t)}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1\mathnormal{:}M],\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}})))$;
6       $\gamma_{\text{+}}^{(t)}\leftarrow[\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{1:n\text{-}1}),\gamma_{\text{+}}^{(t)}\!(\ell_{n}),\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{n\text{+}1:P})]$;
7       foreach $i=1:P$ do
8             Compute $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}})),\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(t)}),\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(t)})$ +++++++++++++++++++++++++++ via Corollary 6;
9
10             $\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))\leftarrow\frac{\widetilde{\pi}_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))}{\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(t)})}$;
11             $\phi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))\leftarrow\alpha\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))$ $\phantom{++++++++++++}+\frac{(1\text{-}\alpha)\widetilde{\pi}_{i}^{\beta}\!(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))}{\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(t)})}$;
12             $\widetilde{\rho}(i|\gamma_{\text{+}}^{(t)})\leftarrow\frac{\phi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}}))}$;
13
14      $\rho(\cdot|\gamma_{\text{+}}^{(t)})\leftarrow\frac{\widetilde{\rho}(\cdot|\gamma_{\text{+}}^{(t)})}{\langle\widetilde{\rho}(\cdot|\gamma_{\text{+}}^{(t)}),1\rangle}$;
15
Algorithm 2 $TGS^{+}$
Input : $\gamma_{\text{+}}^{(0)}$, $\alpha$, $\beta$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$
Output : $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$, $[\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$, $[\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$, $\rho(\cdot|\gamma_{\text{+}}^{(0)})$
1   $P\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\texttt{=}1}^{P},\textsf{\footnotesize row})$;   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\texttt{=}1}^{P},\textsf{\footnotesize col})\texttt{-}2$;
@@
3       $\widetilde{\pi}_{P}(j|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{P}}))\leftarrow 1$;
@@
6       if $\gamma_{\text{+}}^{(0)}\!(\ell_{i})>0$ then
7             $\widetilde{\pi}_{P}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{P}}))\leftarrow 0$;
@@
12             if $\widetilde{\pi}_{P}(j|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{P}}))=0$ $j\neq\gamma_{\text{+}}^{(0)}\!(\ell_{i})$ then
13                   $\widetilde{\pi}_{i}(j|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))\leftarrow 0$;
@@
16                   $\widetilde{\pi}_{i}(j|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))\leftarrow\eta_{i}(j)$;
@@
19      $\pi_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))\leftarrow\frac{\widetilde{\pi}_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))}{\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})}$;
20       $\phi_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))\leftarrow\alpha\pi_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))$ $\phantom{++++++++++++}+\frac{(1\text{-}\alpha)\widetilde{\pi}_{i}^{\beta}\!(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))}{\nu_{i}^{(\beta)}\!(\gamma_{\text{+}}^{(0)})}$;
21       $\widetilde{\rho}(i|\gamma_{\text{+}}^{(0)})\leftarrow\frac{\phi_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}^{(0)}\!(\ell_{i})|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))}$;
@@
The Markov Chain $\{\gamma_{\text{+}}^{(t)}\}_{t\texttt{=}1}^{\infty}$ generated by the TGS kernel (Algorithm 2) is ergodic, and converges to a stationary distribution (not necessarily $\pi$). Additionally, taking into account the importance weight $w^{(t)}\propto P/\left\langle\rho(\cdot|\gamma_{\text{+}}^{(t)}),1\right\rangle$, the weighted samples converge to the stationary distribution $\pi$ in the sense that for any bounded function $h:\Gamma\!_{\text{+}}\rightarrow\mathbb{R}$
$\lim_{T\rightarrow\infty}\sum_{t\texttt{=}1}^{T}w^{(t)}h(\gamma_{\text{+}}^{(t)})=\sum_{\gamma_{\text{+}}\in\Gamma\!_{\text{+}}}h(\gamma_{\text{+}})\pi(\gamma_{\text{+}}).$ (28)
Further, the variance of the importance weights is stable, i.e., does not grow with the number of coordinates.
Proof:
Since the chain is contained in a finite space $\Gamma\!_{\text{+}}$, the convergence results follows from [38, Proposition 1], and the stability of the importance weights follows from [49]. ∎
@@ III-D1 Random-scan GS
  • Sampling $n$ uniformly from $\{1\text{:}P\}$ incurs $\mathcal{O}(P)$ complexity (line 4);
  • Sampling $\gamma_{\text{+}}^{(t)}(\ell_{n})$ from $\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)}\!(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 4);
  • Computing $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$ and $[\nu_{i}^{(1)}\!(\gamma_{\text{+}})]_{i\text{=}1}^{P}$ (based on Corollary 6) takes $\mathcal{O}(P)$ complexity (lines 4-4).
Input : $\gamma_{\text{+}}^{(0)}$, $T$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$, $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(0)}\!(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$, $[\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(0)})]_{i\text{=}1}^{P}$
Output : $[\gamma_{\text{+}}^{(t)}]_{t\text{=}1}^{T}$
1   $P\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize row})$;   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]_{i\text{=}1}^{P},\textsf{\footnotesize col})\texttt{-}2$;
@@ 3       $n\sim\mathcal{U}(1,P)$;
4       $\gamma_{\text{+}}^{(t)}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1\mathnormal{:}M],\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}})))$;
5       $\gamma_{\text{+}}^{(t)}\leftarrow[\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{1:n\text{-}1}),\gamma_{\text{+}}^{(t)}\!(\ell_{n}),\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{n\text{+}1:P})]$;
@@
7             Compute $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(t)}\!(\ell_{\bar{i}})),\nu_{i}^{(1)}\!(\gamma_{\text{+}}^{(t)})$ via Corollary 6;

### v2

@@ III-C Linear Complexity TGS
$TGS^{+}$ also requires the initial coordinate distribution $\rho(\cdot|\gamma_{\text{+}}^{(0)})$ as an input, which can be pre-computed via the initialization routine described in Algorithm 3. The recursive construct (lines 3-3, 3-3) reduces the $\mathcal{O}(P^{2}M)$ complexity [30] to an $\mathcal{O}(PM)$ complexity. Specifically, for each $i\in\{1\mathnormal{:}P\}$, computing $\widetilde{\pi}_{i}(\gamma_{\text{+}}^{(0)}(\ell_{i})|\gamma_{\text{+}}^{(0)}(\ell_{\bar{i}}))$, $\nu_{i}^{(1)}(\gamma_{\text{+}}^{(0)})$, $\nu_{i}^{(\beta)}(\gamma_{\text{+}}^{(0)})$ and $\widetilde{\rho}(i|\gamma_{\text{+}}^{(0)})$ incurs $\mathcal{O}(M)$ complexity.
Input : $\gamma^{(0)}_{\text{+}}$, $\alpha$, $\beta$, $[\eta_{i}(j)]^{P}_{i\text{=}1}$
Output : $[\widetilde{\pi}_{i}(\cdot|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))]^{P}_{i\text{=}1}$, $[\nu^{(1)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$, $[\nu^{(\beta)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$, $\rho(\cdot|\gamma^{(0)}_{\text{+}})$
1   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]^{P}_{i\text{=}1},\textsf{\footnotesize col})\texttt{-}2$;
@@
3     $\widetilde{\pi}_{P}(j|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{P}}))\leftarrow 1$;
@@
6     if $\gamma^{(0)}_{\text{+}}\!(\ell_{i})>0$
7          $\widetilde{\pi}_{P}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{P}}))\leftarrow 0$;
@@
12          if $\widetilde{\pi}_{P}(j|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{P}}))=0$ $j\neq\gamma^{(0)}_{\text{+}}\!(\ell_{i})$
13               $\widetilde{\pi}_{i}(j|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow 0$;
@@
16               $\widetilde{\pi}_{i}(j|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow\eta_{i}(j)$;
@@
19    $\pi_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow\frac{\widetilde{\pi}_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))}{\nu^{(1)}_{i}\!(\gamma^{(0)}_{\text{+}})}$;
20     $\phi_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow\alpha\pi_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))$ $\phantom{++++++++++++}+\frac{(1\text{-}\alpha)\widetilde{\pi}^{\beta}_{i}\!(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))}{\nu^{(\beta)}_{i}\!(\gamma^{(0)}_{\text{+}})}$;
21     $\widetilde{\rho}(i|\gamma^{(0)}_{\text{+}})\leftarrow\frac{\phi_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))}{\pi_{i}(\gamma^{(0)}_{\text{+}}\!(\ell_{i})|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))}$;
@@
Starting from any positive 1-1 initialization, the Markov Chain $\{\gamma_{\text{+}}^{(t)}\}_{t\texttt{=}1}^{\infty}$ generated by the TGS kernel (Algorithm 2) is ergodic, and converges to a stationary distribution (not necessarily $\pi$). Additionally, taking into account the importance weight $w^{(t)}\propto P/\left\langle\rho(\cdot|\gamma_{\text{+}}^{(t)}),1\right\rangle$, the weighted samples converge to the stationary distribution $\pi$ in the sense that for any bounded function $h:\Gamma\!_{\text{+}}\rightarrow\mathbb{R}$
$\lim_{T\rightarrow\infty}\sum_{t\texttt{=}1}^{T}w^{(t)}h(\gamma_{\text{+}}^{(t)})=\sum_{\gamma_{\text{+}}\in\Gamma\!_{\text{+}}}\pi(\gamma_{\text{+}})h(\gamma_{\text{+}}),$ (23)
where the coordinate probability distribution $\rho(\cdot|\gamma_{\text{+}}^{(t)})$ is given by (19). Further, the variance of the importance weights is stable, i.e., does not grow with the number of coordinates.
Proof.
See Appendix (Subsection VI-B). ∎
@@ III-D1 Random-scan GS
Input :  $\gamma^{(0)}_{\text{+}}$, $T$, $[\eta_{i}(j)]^{P}_{i\text{=}1}$, $[\widetilde{\pi}_{i}(\cdot|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))]^{P}_{i\text{=}1}$
Output : $[\gamma^{(t)}_{\text{+}}]^{T}_{t\text{=}1}$
1   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]^{P}_{i\text{=}1},\textsf{\footnotesize col})\texttt{-}2$;
@@ 3     $n\sim\mathcal{U}(1,P)$;
4     $\gamma^{(t)}_{\text{+}}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1\mathnormal{:}M],\widetilde{\pi}_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}})))$;
5     $\gamma^{(t)}_{\text{+}}\leftarrow[\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{1:n\text{-}1}),\gamma^{(t)}_{\text{+}}\!(\ell_{n}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1:P})]$;
@@
7          Compute $\widetilde{\pi}_{i}(\cdot|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))$ via Corollary 1;
@@
Algorithm 4 $RGS^{+}$
  • Sampling $n$ uniformly from $\{1\mathnormal{:}P\}$ incurs $\mathcal{O}(P)$ complexity (line 4);
  • Sampling $\gamma_{\text{+}}^{(t)}(\ell_{n})$ from $\widetilde{\pi}_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}\!(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 4); and
  • Computing $[\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))]_{i\text{=}1}^{P}$ (based on Corollary 1) takes $\mathcal{O}(P)$ complexity (lines 4-4).
III-D2 Deterministic-scan GS
The other extreme of complete randomness, namely deterministic coordinate selection can also be accommodated by $TGS^{+}$. The simplest deterministic scan is to traverse the coordinates according to the periodic sequence $1,2,\cdots,P$, i.e., the selected coordinate at $t$-th iteration is given by $c(t)=1+\text{mod}(t\texttt{-}1,P)$. To realize this in the TGS framework, a proposal of the form
$\varphi_{i}(j|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))=\delta_{c(t)}[i]\phi_{i}(j|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))+\varepsilon(1-\delta_{c(t)}[i]),$
where $\varepsilon$ is a very small positive number. Note that the corresponding coordinate distribution is given by
$\rho(i|\gamma_{\text{+}})\propto\frac{\delta_{c(t)}[i]\phi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))}+\frac{\varepsilon(1-\delta_{c(t)}[i])}{\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))}.$
In principle, this special case has $\mathcal{O}(T(P+M))$ complexity. However, since the coordinate is effectively selected according to the prescribed sequence $c(t)$, the sampling step is not needed. Moreover, using the recursive construct in the initialization (Algorithm 3), it is possible to implement this so-called $DGS^{+}$ special case as described in Algorithm 5, which only incurs an $\mathcal{O}(TM)$ complexity. A similar approach based on the reverse scan order, i.e., $P,P-1,\cdots,1$, is possible by setting $c(t)=P-\text{mod}(t\texttt{-}1,P)$ in Algorithm 5 (line 5).
Input : $\gamma^{(0)}_{\text{+}}$, $T$, $\alpha$, $\beta$, $[\eta_{i}(j)]^{P}_{i\text{=}1}$, $[\widetilde{\pi}_{i}(\cdot|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))]^{P}_{i\text{=}1}$, $[\nu^{(1)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$, $[\nu^{(\beta)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$
Output : $[\gamma^{(t)}_{\text{+}}]^{T}_{t\text{=}1}$
1   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]^{P}_{i\text{=}1},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3     $n\leftarrow c(t)$;   $m\leftarrow c(t-1)$;
4     if $\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{m})>0$
5          $\widetilde{\pi}_{m}(\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{m})|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{m}}))\leftarrow 0$;
6
7     foreach $j=-1:M$ do
8          if $\widetilde{\pi}_{m}(j|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{m}}))=0$ $j\neq\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n})$
9               $\widetilde{\pi}_{n}(j|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))\leftarrow 0$;